OpenStudy (anonymous):

Roll two 6-sided dice and examine their sum. a) How long will it take to roll "snake-eyes" (a pair of ones)? b) What is the probability of rolling a sum of 7 on the first roll? c) What is the probability of rolling a sum of 7 on the fourth roll? d) What is the probability of rolling a sum of 7 by the fourth roll? e) What is the probability of it taking more than 10 rolls to roll the sum of 7?

OpenStudy (kirbykirby):

a) For this one, I think you have to look at it from a geometric distribution point of view. (How many failures will it take to get a success, and your success is the pair of 1's). And the probability of this success (getting two 1's) is 1/36 But, we don't know how many trials are required to do so. So, I think the approach to take (I think I did a similar example once, I'll check my notes when I go home), is to consider: You get it on the 1st try: 1/36 You get it on the 2nd try: 35/36 * 1/36 (you fail the 1st time, but you get it on the 2nd time) 3rd try: 35/36*35/36*1/36 .... until the N'th try. You must then add all these probabilities using an infinite geometric series I believe.

OpenStudy (kirbykirby):

b) 6/36 = 1/6 (You can just enumerate all 6 poissbilities that gives you the sum "7") c) First 3 rolls don't add to 7 with probability 1-6/36 = 30/36 for EACH roll. Thus you have (30/36)^3 * 6/36 d) Some similar argument to a), you consider 1st try: 6/36 2nd try: 30/36 * 6/36 3rd try: (30/36)^2 * 6/36 4th: (30/36)^3 * 6/36 and you just add these probabilities together

OpenStudy (anonymous):

do I add 1st, 2nd,3rd,4th to get the total probabilities

OpenStudy (kirbykirby):

Yes for d)

OpenStudy (anonymous):

ok

OpenStudy (kirbykirby):

for a), I am doubting myself because when I calculated the probability, it just gave 1. But then I realized the question asks "How long it will take", rather than what is the probability... so um I am not too certain for a).

OpenStudy (kirbykirby):

Although I guess you could said that you'd expect to roll about 36 times before get a snake-eyes

OpenStudy (anonymous):

1.35

OpenStudy (kirbykirby):

for e) P(more than 10 rolls to get 7) = P(no 7's on the 1st 10 rolls). We saw that the probability of of getting 7 = 1/6, so the probability of NOT getting 7 in a trial = 1 - 1/6 = 5/6 Hence, P(no 7s in 10 rolls) = (5/6)^10

OpenStudy (kirbykirby):

for d) I got 0.5177

OpenStudy (anonymous):

0.53

OpenStudy (anonymous):

I added 0.17+0.14+0.12+0.10=0.53

OpenStudy (kirbykirby):

oh I added the exact fractions, so this is just rounding error

OpenStudy (anonymous):

ok, I re-added and got 0.53

OpenStudy (anonymous):

so a is 1/36

OpenStudy (kirbykirby):

No I think you should say it takes about 36 rolls to get a snake-eyes. Again, I was thinking of a probability at first, but they ask "How long will it take"

OpenStudy (anonymous):

It looks like d and c is the same question

OpenStudy (anonymous):

I see d say by

OpenStudy (kirbykirby):

They look the same, but they aren't. The "on" and "by" make a difference. in c), it asks the probability ON the 4th roll. This mean, you cannot get a "7" on roll 1, AND can';t get 7 on roll 2, AND can't get 7 on roll 4. Finally, you DO get a 7 on roll 4. In d), it asks BY the 4th roll. This mean's, you can get a 7 on roll 1, OR a 7 on roll 2, OR a 7 on roll 3, OR a 7 on roll 4 (hence, you use the

OpenStudy (kirbykirby):

hence you use the addition rule for each roll probability in d)

OpenStudy (kirbykirby):

in c), it asks the probability ON the 4th roll. This mean, you cannot get a "7" on roll 1, AND can';t get 7 on roll 2, AND can't get 7 on roll 3. ** sorry for the typo.

OpenStudy (anonymous):

That's ok and thank you

OpenStudy (kirbykirby):

:)